# Advancing Identification method of Gamma-Ray Bursts with Data and Feature Enhancement

### v1

@@ 4 Result
All datasets used in this study are derived from Fermi/GBM observational data, divided into three defined subsets for training, validation, and testing. The training set incorporates synthetic GRB samples generated via our proposed data augmentation technique, designed to increase sample size and enhance the diversity of burst signals. The efficacy of this method is visually validated in Figure 1, which illustrates controlled modulation of burst SNR while preserving background levels. Focused on sampling and attenuation factors, the augmentation notably enriches low-SNR GRB representation: as shown in Figure 2, the peak-SNR distribution of GRB samples expands toward lower values, increasing total GRB training samples from 6,189 to over 100,000. Non-GRB samples—sourced from triggerless time-tagged event (TTE) data—were similarly augmented to 100,000, ensuring class balance. For validation and testing, we used 2,774 and 3,143 primary GRBs, respectively, selected based on the number of triggering NaI detectors. Detailed dataset specifications are provided in Table 1. Each input sample consists of light curves across 9 energy bands (64 ms temporal resolution; Figure 3, left), with per-band standardization applied to reduce energy-specific magnitude biases. This preprocessing preserved fine-grained features, enhanced the model’s feature extraction and generalization capabilities, and facilitated subsequent visualization analyses.
@@
Model parameters were optimized via backpropagation and gradient-based methods on our large-scale datasets, enhancing autonomous learning of salient features. Systematic evaluation of hyperparameter configurations identified the optimal set (Table 2). Training and validation loss curves reflect stable dynamics: the early stopping mechanism terminated training at epoch 18 (the optimal point), preserving parameters that generalize best to unseen data. Convergence of both curves—with no divergence between training and validation losses—confirms robust performance without overfitting. We conducted a comparative analysis of multiple models on the test set, including: a baseline ResNet, ResNet variants with existing feature enhancement modules (SE, CBAM, ECA, ASB), and our ResNet-AFFE. All models were trained with hyperparameters consistent with their original literature to ensure fairness, and prediction confidence was estimated via Monte Carlo Dropout (1000 stochastic forward passes). Performance metrics and uncertainties are summarized in Table 3. Incorporating the AFFE module yielded a statistically quantifiable improvement in classification accuracy. Specifically, ResNet-AFFE achieved 97.46% accuracy—outperforming all tested variants with conventional enhancement modules—with consistent results across multiple experimental runs. Grad-CAM visualized the model’s decision-making process, revealing that ResNet-AFFE prioritizes physically meaningful GRB characteristics. As shown in Figure 4, compared to the baseline ResNet, our model’s attention is more strongly concentrated in the interval (containing key prompt emission information) rather than random noise—highlighting a critical difference in feature focus. This visualization enables even untrained users to distinguish model robustness, even for identical predictions.
@@ 5 Discussion and Conclusion
By generating over 100,000 synthetic GRB samples that an order-of-magnitude expansion from the original 6,189 real-world observations, this framework effectively mitigates observational selection bias against low-significance events and resolves the ”observational iceberg” effect described by 2014MNRAS.442.1922L and iceberg-effects_2022Moss, where the majority of astrophysical transients lie below conventional detection thresholds due to instrumental sensitivity limitations and background contamination.
@@
As demonstrated in Figure 2, our methodology successfully reduces the observational bias against low-significance events, which constitute the predominant population of actual GRB phenomena yet remain systematically underrepresented in existing catalogs. Specifically, we bin light curves into 9 distinct energy channels following traditional trigger search algorithms (fermi_gbm_fouth_catalog; search_grb_hxmt_cai), preserving crucial spectral information while enhancing signal detectability. We then systematically modulate burst SNR while retaining realistic background characteristics, yielding physically faithful synthetic samples that span the full dynamic range of actual GRB phenomena. This approach not only reduces model overfitting but also enhances sensitivity to weak transients—critical for capturing the majority of GRBs that lie below conventional detection thresholds due to instrumental constraints. The proposed augmentation framework offers substantial advantages for GRB identification systems by alleviating several fundamental limitations in current approaches. Furthermore, the enhanced sample diversity allows the development of detection algorithms with improved sensitivity to weak transients that typically evade identification in standard analysis pipelines. As such, the augmentation strategy provides a generalizable solution for transient astronomy, where limited mission durations and observational biases often hinder deep learning applications.
@@
We design a novel Adaptive Frequency Feature Enhancement module, AFFE, which represents a methodological advance in addressing the spectral-temporal complexity of GRB signals. Unlike conventional time-domain attention mechanisms (e.g., SE, CBAM, ECA) that fail to capture frequency-domain correlations, or fixed-frequency filters lacking adaptability, AFFE employs a dual-path architecture to perform data-driven spectral weighting—enabling adaptive soft-threshold denoising and coherent feature extraction without manual filter design. This mechanism aligns with the complex, variable frequency signatures of GRBs, optimizing the model’s ability to distinguish GRBs from non-GRB signals by leveraging discriminative frequency components. Integrating AFFE with a ResNet baseline yields a classification accuracy of 97.46%, representing a statistically significant 3% improvement over the state-of-the-art 2D ResNet-CBAM architecture (94.46%; (DL_identify_grb_by_peng)). Comparative analysis of these models and their predictive uncertainty reveal that our AFFE implementation shows superior performance metrics, establishing a new benchmark for high-precision GRB identification. Comparative analysis further validates that this frequency-adaptive architecture maintains performance advantages even near theoretical accuracy saturation, demonstrating its superiority and robustness in handling the unique spectral dynamics of GRB light curves.
@@
Further validation confirms that this frequency-adaptive architecture maintains performance advantages even near theoretical accuracy saturation, demonstrating its superiority and robustness in handling the unique spectral dynamics of GRB light curves, highlighting the methodological advancement in alleviating the unique challenges of GRB identification. As visualized in Figure 4, the model selectively attends to critical burst characteristics, exhibiting strong concordance with domain expertise in GRB identification. The incorporation of the AFFE module significantly enhances the ResNet architecture’s capacity to extract physically meaningful features, enabling comprehensive characterization of burst processes. This Fourier-domain enhancement improves classification by explicitly modeling spectral-temporal correlations—critical for distinguishing GRBs from noise, as key burst signatures (e.g., multi-phase emission, spectral lags) are jointly encoded in both time and frequency domains. It is important to note that predictive uncertainty remains a critical consideration in real-world astrophysical applications. while our model achieves high accuracy, the confidence of predictions—especially for low-significance or atypical bursts—requires further quantification, and future work should integrate uncertainty estimation techniques to better assess classification reliability.
@@
We also gain valuable physical insights from model-derived features. The dimensionality reduction and AutoGMM clustering of model-extracted features uncover meaningful physical patterns in GRB populations. Kilonova-associated GRBs (GRB-KN) predominantly cluster in Clusters 2 and 3, with long-duration GRB-KN in Cluster 2 sharing multi-phase light curve morphologies—supporting a common progenitor system as proposed by previous theoretical work (KN-GRB_Zhu2022ApJ), suggesting a common progenitor system for these apparently long-duration events. A particularly compelling case is presented by GRB 211211A and GRB 230307A, which exhibit nearly identical three-phase emission structures (precursor, main burst, extended emission) and occupy adjacent regions in the low-dimensional embedding space (DownDimGRB_Dimple2023; DownDimGRB_Negro2025). The comprehensive temporal and spectral analyses of their strikingly similar properties indicate a common mechanism (GRB-KN-Compare_2024Peng; GRB-KN-Compare_2025Wang). This clustering pattern mirrors yet remains distinct from the tight grouping of GRB-SN in Clusters 1 and 2, which similarly reflects their shared progenitor physics (GRB-SN-similar_2024Kumar; GRB-SN-similar_2025Kumar). GRBs with extended emission show spatial separation into three subgroups highly correlated with duration, while ultra-long GRBs are confined to the lGRB-dominated Cluster 3—hinting at potential subclasses with physical origins independent of standard duration-based classification. Notably, sGRBs with precursors occupy transitional zones between short and long GRB clusters, indicating hybrid characteristics and diverse progenitor systems. In contrast, lGRBs with precursors show widespread dispersion, implying precursor activity is an independent physical process rather than a core burst characteristic. These findings reinforce the limitations of duration-only GRB classification, aligning with prior studies (e.g., (GRB-cluter_Modak_2021)) that highlight inherent uncertainties in this paradigm. The model’s ability to aggregate GRBs by physical origin rather than just observational parameters demonstrates that machine learning-derived features can serve as a complementary dimension for probing GRB progenitors, with future increases in the sample size of events with confirmed origins essential for further validating these classification schemes.
@@
Despite achieving high classification accuracy, the model exhibits residual misclassification of low-SNR GRBs (99% of false negatives have peak-SNR 5 ), which may require specialized feature engineering or multi-modal data integration (e.g., combining light curves with spectral data, energy response-corrected data) for further improvement. Predictive uncertainty quantification—critical for real-world astrophysical applications—also warrants deeper integration, as confidence assessment for low-significance or atypical bursts remains underdeveloped. Additionally, validating the clustering results for special GRB subclasses (e.g., GRB-KN, precursor-containing GRBs) requires larger samples of events with confirmed origins. Future work will focus on three key directions: (1) integrating Monte Carlo Dropout or Bayesian neural networks to quantify prediction uncertainty and improve reliability for faint events; (2) expanding the framework to process multi-modal data (spectral, temporal, and polarization information) for richer feature extraction; (3) adapting the model for real-time analysis pipelines in time-domain and multi-messenger astronomy, enabling faster detection and characterization of transient phenomena. The adaptive frequency-domain approach and data augmentation strategy developed here are also generalizable to other transient astrophysical events (e.g., fast radio bursts, soft gamma-ray repeaters, supernovae), highlighting their broader potential for advancing observational capabilities.
@@ 6 Acknowledgements
We would like to thank Prof Rui Luo, Dr. Yi Yang, and Prof Jia-wei Luo for helpful discussion. This study is supported by the National Key R&D Program of China (2024YFA1611703), the National Natural Science Foundation of China (grant Nos. 12473044, 12494572, 12273042, 12133007, 41827807 and 61271351) and the Science and Technology Innovation Plan of Shanghai Science and Technology Commission (22DZ1209500). This work is also partially supported by the Strategic Priority Research Program of the CAS under grant No. XDA15360300. B. L. acknowledges support from the National Astronomical Science Data Center Young Data Scientist Program (grant No. NADC2023YDS-04).

### v2

@@ 4 Result
All datasets used in this study are derived from Fermi/GBM observational data, divided into three defined subsets for training, validation, and testing. The training set incorporates synthetic GRB samples generated via our proposed data augmentation technique, designed to increase sample size and enhance the diversity of burst signals. The efficacy of this method is visually validated in Figure 1, which illustrates controlled modulation of burst SNR while preserving background levels. Focused on sampling and attenuation factors, the augmentation notably enriches low-SNR GRB representation: as shown in Figure 2, the peak-SNR distribution of GRB samples expands toward lower values, increasing total GRB training samples from 6,189 to over 100,000. Non-GRB samples, sourced from triggerless time-tagged event data, were similarly augmented to 100,000, ensuring class balance. For validation and testing, we used 2,774 and 3,143 primary GRBs, respectively, selected based on the number of triggering NaI detectors. Detailed dataset specifications are provided in Table 1. Each input sample consists of light curves across 9 energy bands (64 ms temporal resolution; Figure 3, left), with per-band standardization applied to reduce energy-specific magnitude biases. This preprocessing preserves fine-grained features, enhanced the model’s feature extraction and generalization capabilities, and facilitated subsequent visualization analyses.
@@
Model parameters were optimized via backpropagation and gradient-based methods on our large-scale datasets, enhancing autonomous learning of salient features. Systematic evaluation of hyperparameter configurations identified the optimal set (Table 2). Training and validation loss curves exhibit stable dynamics: the early stopping mechanism terminated training at epoch 18 (the optimal point), preserving parameters that generalize best to unseen data. Convergence of both curves—with no divergence between training and validation losses—confirms robust performance without overfitting. We conducted a comparative analysis of multiple models on the test set, including: a baseline ResNet, ResNet variants with existing feature enhancement modules (SE, CBAM, ECA, ASB), and our ResNet-AFFE. All models were trained with hyperparameters consistent with their original literature to ensure fairness, and prediction confidence was estimated via Monte Carlo Dropout (1000 stochastic forward passes). Performance metrics and uncertainties are summarized in Table 3. Incorporating the AFFE module yielded a statistically quantifiable improvement in classification accuracy. Specifically, ResNet-AFFE achieved 97.46% accuracy—outperforming all tested variants with conventional enhancement modules—with consistent results across multiple experimental runs. Grad-CAM visualized the model’s decision-making process, revealing that ResNet-AFFE prioritizes physically meaningful GRB characteristics. As shown in Figure 4, compared to the baseline ResNet, our model’s attention is more strongly concentrated in the interval (containing key prompt emission information) rather than random noise—highlighting a critical difference in feature focus. This visualization enables even untrained users to distinguish model robustness, even for identical predictions.
@@ 5 Discussion and Conclusion
By generating over 100,000 synthetic GRB samples that an order-of-magnitude expansion from the original 6,189 real-world observations, this framework effectively mitigates observational selection bias against low-significance events and helps to address the ”observational iceberg” effect described by Lü et al. (2014) and Moss et al. (2022), where the majority of astrophysical transients lie below conventional detection thresholds due to instrumental sensitivity limitations and background contamination.
@@
As demonstrated in Figure 2, our data augmentation methodology successfully reduces the observational bias against low-significance events, which constitute the predominant population of actual GRB phenomena yet remain systematically underrepresented in existing catalogs. Specifically, we bin light curves into 9 distinct energy channels following traditional trigger search algorithms (von Kienlin et al., 2020; Cai et al., 2021), preserving crucial spectral information while enhancing signal detectability. We then systematically modulate burst SNR while retaining realistic background characteristics, yielding physically faithful synthetic samples that span the full dynamic range of actual GRB phenomena. This approach not only reduces model overfitting but also enhances sensitivity to weak transients—critical for capturing the majority of GRBs that lie below conventional detection thresholds due to instrumental constraints. The proposed augmentation framework offers substantial advantages for GRB identification systems by alleviating several fundamental limitations in current approaches. Furthermore, the enhanced sample diversity allows the development of detection algorithms with improved sensitivity to weak transients that typically evade identification in standard analysis pipelines. As such, the augmentation strategy provides a generalizable solution for transient astronomy, where limited mission durations and observational biases often hinder deep learning applications.
@@
We design a novel Adaptive Frequency Feature Enhancement module, AFFE, which represents a methodological advance in addressing the spectral-temporal complexity of GRB signals. Unlike conventional time-domain attention mechanisms (e.g., SE, CBAM, ECA) that fail to capture frequency-domain correlations, or fixed-frequency filters lacking adaptability, AFFE employs a dual-path architecture to perform data-driven spectral weighting—enabling adaptive soft-threshold denoising and coherent feature extraction without manual filter design. This mechanism aligns with the complex, variable frequency signatures of GRBs, optimizing the model’s ability to distinguish GRBs from non-GRB signals by leveraging discriminative frequency components. Integrating AFFE with a ResNet baseline yields a classification accuracy of 97.46%, representing a statistically significant 3% improvement over the state-of-the-art 2D ResNet-CBAM architecture (94.46%; (Zhang et al., 2024)). Comparative analysis of these models and their predictive uncertainty reveal that our AFFE implementation shows superior performance metrics, establishing a new benchmark for high-precision GRB identification. Comparative analysis further validates that this frequency-adaptive architecture maintains performance advantages even near theoretical accuracy saturation, demonstrating its superiority and robustness in handling the unique spectral dynamics of GRB light curves.
@@
Further validation confirms that this frequency-adaptive architecture maintains performance advantages even near theoretical accuracy saturation, demonstrating its superiority and robustness in handling the unique spectral dynamics of GRB light curves, highlighting the methodological advancement in alleviating the unique challenges of GRB identification. As visualized in Figure 4, the model selectively attends to critical burst characteristics, exhibiting strong concordance with domain expertise in GRB identification. The incorporation of the AFFE module significantly enhances the ResNet architecture’s capacity to extract physically meaningful features, enabling comprehensive characterization of burst processes. This Fourier-domain enhancement improves classification by explicitly modeling spectral-temporal correlations—critical for distinguishing GRBs from noise, as key burst signatures (e.g., multi-phase emission, spectral lags) are jointly encoded in both time and frequency domains.
@@
We also gain valuable physical insights from model-derived features. The dimensionality reduction and AutoGMM clustering of model-extracted features uncover meaningful physical patterns in GRB populations. Kilonova-associated GRBs (GRB-KN) predominantly cluster in Clusters 2 and 3, with long-duration GRB-KN in Cluster 2 sharing multi-phase light curve morphologies—supporting a common progenitor system as proposed by previous theoretical work (Zhu et al., 2022), suggesting a common progenitor system for these apparently long-duration events. A particularly compelling case is presented by GRB 211211A and GRB 230307A, which exhibit nearly identical three-phase emission structures (precursor, main burst, extended emission) and occupy adjacent regions in the low-dimensional embedding space (Dimple et al., 2023; Negro et al., 2025). The comprehensive temporal and spectral analyses of their strikingly similar properties indicate a common mechanism (Peng et al., 2024; Wang et al., 2025). This clustering pattern mirrors yet remains distinct from the tight grouping of GRB-SN in Clusters 1 and 2, which similarly reflects their shared progenitor physics (Kumar & Sharma, 2024; Kumar, 2025). GRBs with extended emission show spatial separation into three subgroups highly correlated with duration, while ultra-long GRBs are confined to the lGRB-dominated Cluster 3—hinting at potential subclasses with physical origins independent of standard duration-based classification. Notably, sGRBs with precursors occupy transitional zones between short and long GRB clusters, indicating hybrid characteristics and diverse progenitor systems. In contrast, lGRBs with precursors show widespread dispersion, implying precursor activity is an independent physical process rather than a core burst characteristic. These findings reinforce the limitations of duration-only GRB classification, aligning with prior studies (e.g., (Modak, 2021)) that highlight inherent uncertainties in this paradigm. The model’s ability to aggregate GRBs by physical origin rather than just observational parameters demonstrates that machine learning-derived features can serve as a complementary dimension for probing GRB progenitors, with future increases in the sample size of events with confirmed origins essential for further validating these classification schemes.
@@
Despite achieving high classification accuracy, the model exhibits residual misclassification of low-SNR GRBs (99% of false negatives have peak-SNR 5 ), which may require specialized feature engineering or multi-modal data integration (e.g., combining light curves with spectral data, energy response-corrected data) for further improvement. Predictive uncertainty quantification—critical for real-world astrophysical applications—also warrants deeper integration, as confidence assessment for low-significance or atypical bursts remains underdeveloped. Additionally, validating the clustering results for special GRB subclasses (e.g., GRB-KN, precursor-containing GRBs) requires larger samples of events with confirmed origins. Future work will focus on three key directions: (1) developing robust, real-time uncertainty estimation to deliver reliable confidence assessments for faint and atypical transients, which is critical for triggering follow-up observations. (2) expanding the framework to process multi-modal data (spectral, temporal, and polarization information) for richer feature extraction; (3) adapting the model for real-time analysis pipelines in time-domain and multi-messenger astronomy, enabling faster detection and characterization of transient phenomena. The adaptive frequency-domain approach and data augmentation strategy developed here are also generalizable to other transient astrophysical events (e.g., fast radio bursts, soft gamma-ray repeaters, supernovae), highlighting their broader potential for advancing observational capabilities.
@@ 6 Acknowledgements
We would like to thank Prof Rui Luo, Dr. Yi Yang, Prof Jia-Wei Luo, and Rui-Ze Shi for helpful discussion. This study is supported by the National Key R&D Program of China (2024YFA1611703 and 2024YFA1611700), the National Natural Science Foundation of China (grant Nos. 12473044, 12494572, 12273042, 12133007, 41827807 and 61271351) and the Science and Technology Innovation Plan of Shanghai Science and Technology Commission (22DZ1209500). This work is also partially supported by the Strategic Priority Research Program of the CAS under grant No. XDA15360300. B. L. acknowledges support from the National Astronomical Science Data Center Young Data Scientist Program (grant No. NADC2023YDS-04).